SUBMILLIMETER-WAVE SPECTROSCOPY OF THE VAN DER WAALS BENDING BAND OF Ar-HBr

\underline{KENSUKE HARADA}, ASAO MIZOGUCHI, STEPHANE BAILLEUX, and KEIICHI TANAKA, Institute for Molecular Science, Myodaiji, Okazaki, 444-8585 Japan.

The Ar-HBr cluster has a large amplitude intermolecular bending motion. The excited states of the van der Waals vibrations were observed by infrared^{a} and far infrared^{b} laser spectroscopy. The cluster has a nearly ”linear” structure of Ar-HBr in the ground state (Σ_{0} state), while the cluster is estimated to have a nearly ”anti-linear” structure of Ar-BrH in the first excited state (Σ_{1} state) of the intermolecular bending vibration.

In the present study, we have observed submillimeter-wave transitions of the Σ_{1} - Σ_{0} band of Ar-HBr generated in a pulsed supersonic jet expansion. The observed band origins of the Σ_{1} - Σ_{0} bands of Ar-H^{79}Br and Ar-H^{81}Br are 6 and 27 MHz lower than those reported by combination differences of infrared data^{a}. The eQq constants in the Σ_{1} states were determined for the first time to be 260.90(12) and 217.854(98) MHz for Ar-H^{79}Br and Ar-H^{81}Br, respectively, which agree well with the estimated values (238 and 199 MHz for Ar-H^{79}Br and Ar-H^{81}Br) from a potential calculation^{c}. The cos^{−1}\sqrt{⟨cos^{2}θ⟩} value is 144.3^{∘} in the Σ_{1} state, which is different from the value, 42.1^{∘}, in the ground state, where the θ is an angle between a cluster axis and a HBr monomer axis.

aJ. Han, A. L. McIntosh, Z. Wang, R. R. Lucchese, and J. W. Bevan, Chem. Phys. Lett. 265, 209 (1997).

bD. W. Firth, M. A. Dvorak, S. W. Reeve, R. S. Ford, and K. R. Leopold, Chem. Phys. Lett. 168, 161 (1990).

cJ. M. Hutson, J. Chem. Phys. 91, 4455 (1989).